You manage a risky portfolio with an expected rate of return of 19% and a standard deviation of 33%. The T-bill rate is 7%. Your client chooses to invest 80% of a portfolio in your fund and 20% in a T-bill money market fund.

What is the reward-to-volatility (Sharpe) ratio (S) of your risky portfolio? Your client’s? (Do not round intermediate calculations. Round your answers to 4 decimal places.)

Your reward-to-volatility ratio?________

Clients' reward-to-volatility ratio?_________
Answer #1

reward to volatility ratio = (Return - riskfree rate)/standard deviation

your reward to volatility ratio = (19%-7%)/33%

= 0.3636

client return = 0.8 * 19% + 0.2 * 7% = 16.6%

client standard deviation = 0.8 * 33% = 26.4

client reward to volatility ratio = (16.6%-7%)/26.4%

= 0.3636